# LLMs’ Leaning in European Elections

### v1

@@ Abstract
Many studies suggest that LLMs have left wing leans [1, 2, 3, 4, 5, 6]. The article extends the US presidential election analysis made in [5], where multiple LLMs were asked to vote between Joe Biden and Donald Trump in a virtual election, and the results showed a clear lean of LLMs toward Joe Biden. This article considers natural follow-up questions that could arise from that experiment, such as: what is the extent of this phenomenon? Is it generalizable to multiple virtual elections in other countries? How can we lean the models toward a different choice?
@@ 3 Methodology
To extract preference information from LLMs, we used a forced choice format [20, 21]. We ask multiple LLMs to answer a very simple question: between two political candidates, one from the left, and one from the right wing, what is the candidate that would you vote for? We asked the same question in two different experiments. In the first one, we did not ask the models to impersonate a specific type of person, in the second one we asked the model to impersonate two different types of persona. For the two experiments, we used the same inference parameters, reported in Table  2. These parameters were selected to not modify the original token distribution of the models. Each experiment is repeated 100 times and the order between the candidates is switched 50% of the times to remove possible positional biases of the models. The output of the models has been automatically parsed to extract their votes, and this automatic phase is followed by a manual inspection. Sometimes, models refused to respond to the prompts used in the two experiments, as they were trained to avoid answering political questions. We report all the results, specifying the number of times the models refused to answer.
Table 2: Parameters
Temperature Top-P
1.0 1.0
@@
Given the possibility to vote between two candidates, one from the left wing and the other from the right wing, what is the preferred choice? We considered elections in several European countries, Table  3. We asked each LLM to vote for one of the two candidates. The prompt used for this experiment is the Unbiased Person prompt. We tried to keep the prompt very simple to avoid the introduction of possible biases in the question, but we are aware that asking an LLM to vote for a candidate, is not free from steering effects on its behavior.

### v2

@@ Abstract
Many studies suggest that LLMs lean towards the left [1, 2, 3, 4, 5, 6]. The article extends the analysis of US presidential elections made in [5] considering several virtual elections in multiple European countries. The analysis considers multiple LLMs and the results confirm the extent of the leaning. Furthermore, the results show that the leaning is not uniform between countries. Sometimes, models refuse to take a position in the virtual elections, but the refusal rate itself is not uniform between countries.
@@ 3 Methodology
To extract preference information from LLMs, we used a forced choice format [20, 21]. We ask multiple LLMs to answer a very simple question: between two political candidates, one from the left, and one from the right wing, what is the candidate that would you vote for? We asked the same question in two different experiments. In the first one, we did not ask the models to impersonate a specific type of person, in the second one we asked the model to impersonate two different types of persona. For the two experiments, we used the same inference parameters: temperature=1.0 and top-p=1.0. These parameters were selected to not modify the original token distribution of the models. Each experiment is repeated 100 times and the order between the candidates is switched 50% of the times to remove possible positional biases of the models. The output of the models has been automatically parsed to extract their votes, and this automatic phase is followed by a manual inspection. Sometimes, models refused to respond to the prompts used in the two experiments, as they were trained to avoid answering political questions. We report all the results, specifying the number of times the models refused to answer.
@@
Given the possibility to vote between two candidates, one from the left wing and the other from the right wing, what is the preferred choice? We considered elections in several European countries, Table  2. We asked each LLM to vote for one of the two candidates. The prompt used for this experiment is the Unbiased Person prompt. We tried to keep the prompt very simple to avoid the introduction of possible biases in the question, but we are aware that asking an LLM to vote for a candidate, is not free from steering effects on its behavior.